OpenStudy (anonymous):

Solve the equation. Check the solution. (6r -7) divided by 10 = r divided by 4.

OpenStudy (anonymous):

\[\frac{ 6r - 7 }{ 10 } = \frac{ r }{ 4 }\]

OpenStudy (anonymous):

you mean 6r-7 mod 10=r mod 4? you have to explain a bit..

OpenStudy (anonymous):

I put the equation up

OpenStudy (anonymous):

i thought it was more difficult than this :P

OpenStudy (anonymous):

I can explain the steps

OpenStudy (anonymous):

Unless you want to?

OpenStudy (anonymous):

\[4(6r - 7) = 10r\]

OpenStudy (anonymous):

Cross multiplication

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

\[24 r - 28 = 10 r\]

OpenStudy (anonymous):

Isolate the r on one side and the number on the other

OpenStudy (anonymous):

just write it all at once :P

OpenStudy (anonymous):

\[14r = 28\]

OpenStudy (anonymous):

r = 2

OpenStudy (anonymous):

i can't its a two sided equation xD

OpenStudy (anonymous):

Thank you! :)!!!